OpenStudy (anonymous):

If the length of a rectangle is represented by 4n+5 and the width is represented by 3n-1, which algebraic expression represents the perimeter of the rectangle? a) 4n + 8 b) 7n + 4 c) 7n^2 + 11n - 5 d) (1/2)(7n +4)

OpenStudy (zpupster):

2l + 2W = perimeter 2(4n+5) + 2(3n-1) = P 8n + 10 + 6n - 2 = P 14n + 8 = P looks a and a typo

OpenStudy (anonymous):

thats what i thought, just wanted to make sure